OpenStudy (anonymous):

(a) Factorize a^(3)-8b^(3) and a^(3)-4ab^(2) into linear or quadratic factors with integral coefficients. (b) Find the LCM of a^(3)-8b^(3) and a^(3)-4ab^(2).

OpenStudy (anonymous):

@Callisto would you kindly help me?

OpenStudy (anonymous):

What kind of sorcery is this.

OpenStudy (callisto):

For part a: a^(3)-8b^(3) = a^3 - (2b)^3 Use the identity a^3 - b^3 = (a-b)(a^2 + ab +b^2) to factorize it. a^(3)-4ab^(3) Take out the common factor first. Then, you need to use the identity a^2 - b^2 = (a+b)(a-b) to factorize it.